CRY+VIS Interface Meeting
Date: 2022/1/28 13:00-13:40
Participants: Takahashi, Ushiba, Tomaru, Hirata, Yasui, Ikeda, Uchiyama, Node, Aso, Yokozawa, Akutsu
- IX: Health check was almost finished. Glitch in the F3 signals. Windshield for the MN oplev.
- EX: Health check will be done after the vacuum issue. Leake check on 1/25. Asked the KEK staff to repair the feedthrough in Feb.
- IY: Height adjustment was scheduled on 2/1.
- EY: Height adjustment on 1/25. Oplev setting in Feb. Preparation for the accelerometers. Height measurement was scheduled on 2/1.
- Replacement of flip cables: Type-C.
- TM height setting: ITMY.
- Oplev: Setup of the TM Oplev in IY after the TM alignment in late Feb.
- ACC: Preparation for EY.
- Commissioning in air: Return of GRY. Fluctuation of ETMY is large. MN oplev control is necessary.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: EXV (transportation on 2/7, implementation on 2/8-11)
Heater control: Test in IXV/EXV is planed (Sequence). Power controller in IY.
- Power source for Oplev: IY.
- Tunnel for the IYC booth: Scheduled on 1/31.
- TMS: Scheduled on 2/14-16.
Date: 2022/1/28 14:00-14:25
Participants: Takahashi, Ushiba, Fabian, Hirata, Aso, Tanaka, Terrence, Uchiyama, Sato, Ishizaki
- Adjusted the height of ETMY. The F1 damper was set in a gap of 2mm.
- Repaired the actuator coil of ACC #3 in EY. Investigated the noise level of the accelerometers in IY.
- [Terrence] reworked ETMY controls after the height adjustment.
- F0-F2 transfer functions changed drastically so I reworked the controls.
Tried to offload most DoFs but F0 yaw stepper cannot bring BF yaw to zero.
- Found a 67 mHz very high Q mode that cannot be damped using MN2TM DC control.
- I reduced the UGF of the MN2TM DC control for yaw to 10 mHz and added a notch at 67 mHz so it doesn't interact with the hum.
- MN2TM DC control can be engaged when the oscillation falls well within the oplev range. But the oscillation is damping very very slowly.
- I haven't tried but auto alignment with Guardian may not work because of this oscillation.
- Terrence reported the BS oplev was out of range, so we recovered the yaw alignment using the F0 yaw stepper motor. We had to move the length sensing QPD but we took note of the initial and final values in case we want to go back. Terrence measured IM TFs to make sure the payload didn't touch the security structure.
- We assessed the height of the BS:
- We set BF and F0 to zero
- We chose the position of F1 so the optic was at the desired height.
- The amount of actuation F1 required was about -24,000 counts, which is very large, and the keystone moved around -700 um, which is a very small amount given the tools we're using to measure height.
- Given that the reproducibility of the measurement is within 700 um or so (e. g. PRM SF), for the sake of having a low actuation, we could choose a setpoint higher than the one we measured this week.
- We calibrated the BS F1 LVDT using the optical displacement sensor.
- We used a modified holder that allowed us to adjust an additional DoF.
- We realized this device only works with specular reflection, so we had to tilt the sensor a lot (per the manual) and put think ballast mass with a more reflective surface.
- We moved the F1 keystone with coil-magnet actuation only and the range was small. We should use the fishing rod also.
- We got an straight line, which is very good news.
- We should calibrate again with a larger displacement range.
- Terrence made controls for BS and Guardian can bring it from SAFE and ALIGNED stage.
- There's a 46 mHz hum seen in BS yaw, particularly in IM yaw and TM yaw. TM DC control can suppress the motion at TM to a large extent but the hum is persistent.
We fixed a traverser cable in PR3. An in-vacuum connector was failing and Hirata-san had to crimp the pins again.
- Type-A team: Height adjustment of ITMY, Accelerometers in EY.
- Type-B team: PR2 cable (BF FR, F0 FR), BS SF calibration, PR cable clamp.
(Date in Kamioka)
The week from 1/31
- Takahashi: 2/1-4
- Sato: 2/1-4
- Hirata: 2/1-4
The week from 2/7
- Takahashi: 2/8-10
- Sato: 2/8-10
- Hirata: 2/8-10